OpenStudy (anonymous):

@texaschic101 Can you check 5 more

OpenStudy (texaschic101):

what ya got

OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 5 2x - 4y = 10

OpenStudy (anonymous):

I got more than one solution...

OpenStudy (texaschic101):

x - 2y = 5 2x - 4y = 10 --> reduce x - 2y = 5 same line....infinite solutions

OpenStudy (anonymous):

I got many solutions but there's no option.

OpenStudy (anonymous):

A pair of linear equations is shown below: y = –2x + 3 y = –4x – 1 Which of the following statements best explains the steps to solve the pair of equations graphically? Graph the first equation which has slope = 3 and y-intercept = –2, graph the second equation which has slope = –1 and y-intercept = –4, and find the point of intersection of the two lines. Graph the first equation which has slope = –3 and y-intercept = 2, graph the second equation which has slope = 1 and y-intercept = 4 and find the point of intersection of the two lines. Graph the first equation which has slope = –2 and y-intercept = 3, graph the second equation which has slope = –4 and y-intercept = –1, and find the point of intersection of the two lines. Graph the first equation which has slope = 2 and y-intercept = –3, graph the second equation which has slope = 4 and y-intercept = 1, and find the point of intersection of the two lines.

OpenStudy (anonymous):

A system of equations is shown below: 5x + 2y = 3 (equation 1) 2x – 3y = 1 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof? Show that the solution to the system of equations 7x – y = 4 and 2x – 3y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 2x + 5y = 3 and 3x – 2y = 1 is the same as the solution to the given system of equations Show that the solution to the system of equations 9x + 4y = 5 and 7x – y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations –4x + 9y = 5 and 2x – 3y = 1 is the same as the solution to the given system of equations

OpenStudy (anonymous):

2 different questions

OpenStudy (anonymous):

For the first Graph the first equation which has slope = –3 and y intercept = 2; graph the second equation which has slope = 1 and y intercept = 4 and find the point of intersection of the two lines.

OpenStudy (texaschic101):

graph 1st eq which has slope of -2 and y int of 3 and graph 2nd eq which has slope of - 4 amd y int of - 1......(C) ================== thats the first one...because in y = mx + b form, the slope is in the m position and the y intercept is in the b position

OpenStudy (texaschic101):

I am so sorry...I lost track of time...I have to go....SORRY...I will be back on soon

OpenStudy (anonymous):

its okay thank you